OpenStudy (anonymous):

factorise the polynomial 1000- 27x^6

OpenStudy (anonymous):

http://www.ehow.com/how_4499707_factor-cubes.html

OpenStudy (anonymous):

didnt help sorry

OpenStudy (ajprincess):

1000=10^3 27x^6=3^3(x^2)^3 1000-27x^6=10^3-(3x^2)^3 this is of the form (a^3-b^3)=(a-b)(a^2+b^2+ab) Instead of a apply 10 and b apply 3x^2 Then you will get (10-3x^2)(100+9x^4+30x^2) Hope that helps

OpenStudy (anonymous):

omg that's exaactly wat i was looking 4! thanx loads. this helped a lot! :)

OpenStudy (ajprincess):

you are welcome.:)